# High-Order Hermite Optimization: Fast and Exact Gradient Computation in Open-Loop Quantum Optimal Control using a Discrete Adjoint Approach

### v2

@@ 4.2.2 Performing Gate Optimization
We now perform the gate optimization. We compare the performance of each order of the Hermite method for target relative final-time errors of $10^{-1}$, $10^{-3}$, $10^{-5}$, and $10^{-7}$. The optimization is managed by IPOPT (Interior Point OPTimizer) [39], a software library for solving large scale nonlinear optimization problems using a primal-dual interior point method. We configure IPOPT to perform a limited-memory quasi-Newton approximation of the Hessian of the objective function for use in line searches (second-order optimization). IPOPT also allows for box constraints on the control vector, which we use to enforce our constraint on the control pulse amplitudes.
@@
For each order and target error, we perform the optimization from ten initial control vectors, sampled from a uniform random distribution with a range of $[-0.005,0.005]$. The upper and lower bounds on each control parameter (enforced by IPOPT) are set to ±$\pm 0.05$, enforcing a maximum amplitude of 50 MHz on the B-spline envelope of each carrier wave for each control pulse, the same as the upper and lower limits on the random distribution in the stepsize experiment. The same ten initial control vectors are used for all orders and all target errors, so that every optimization starts from the same analytic control pulses. Each optimization is allowed to run for six hours before being halted.
@@
In Figure 4, for each target error we plot the generalized gate infidelity against the number of IPOPT iterations completed for each order of the Hermite method. In the figure, we do not plot the “watchdog” iterations, in which the optimizer aggressively explores new search directions, resulting in large changes in the objective function which would make the graph difficult to interpret (but the watchdog iterations are still counted in the total). These samples were collected using the same setup as the stepsize test, except we use four CPU cores instead of one and use multithreading to parallelize the forward and adjoint evolutions over the different initial and terminal conditions. In Figure 5, we plot the time evolution of the population of the computational basis states for the control found that best minimized the generalized gate infidelity for a target final state error of $10^{-7}$.
@@
Figure 4: Optimization of a CNOT gate using two qudits coupled to a resonator bus as the mode, using B-spline curves multiplied carrier waves as the control pulse ansatz, using the methods and numbers of timesteps prescribed by Table 6 in order to reach target relative final state errors on the orders of $10^{-1}$, $10^{-3}$, $10^{-5}$, and $10^{-7}$.
@@
Figure 5: Time evolution of the state populations for the CNOT gate by the best (lowest generalized infidelity) control pulses found by the optimizations performed with a target relative final state error of $10^{-7}$ for the two qudits plus resonator model (see Figure 4).
@@ 4.2.3 Analysis
From the the plot of the number of the mean relative error against the number of timesteps for the three subsystem example in Figure 3, we see that each order of the Hermite exhibits the full order of convergence. From the plot of the mean gradient computation time against the number of timesteps, we can see that the computation time increases with the number of timesteps, and the asymptotic behavior of each order of the method is the same, indicating that for the computation time of the higher-order methods is roughly a constant multiple of the computation time of the lower-orders, which is to be expected. Interestingly, for the sixth-twelfth-order methods the computation time decreases around $2^{9}-2^{11}$ timesteps, for which the stepsize ranges from $1.07\geq\Delta t\geq 0.27$ nanoseconds. We attribute this to a decrease in the number of GMRES iterations required in the linear solves of the forward and adjoint evolutions due to the effectiveness of our preconditioner and our initial guess (using Taylor expansion) when $\Delta t<1$ nanosecond.
@@
The computational efficiency of the high-order methods is showcased by Tables 7 and 9. When using the initial conditions from the gate design problem and trying to reach final state errors of $10^{-3}$-$10^{-5}$, using high-order methods results in average speedup factors of approximately $3$-$13$ compared to Störmer-Verlet (when choosing the most efficient order of the method for each target error). When the initial condition is the highly excited state $|933\rangle$, the average speedup factors are in the range $27$-$143$!
In addition to prescribing the number of timesteps required to reach a target error, Tables 6 and 8 also highlight another advantage of the high-order methods: reduced memory requirements. The memory required to store the state history in the forward and adjoint evolutions scales linearly in the number of timesteps and the dimension of Hilbert space. Because the high-order methods can obtain highly accurate solutions with a small number of timesteps, we are often able to store the entire state history in memory. This allows us to avoid using the “memory-lean” version of the gradient calculation, which would require more computation and produce less accurate gradients.
Figure 4 highlights the effectiveness of high-order methods for the optimization process. We observe that for different order methods and different target errors, but starting from the same initial control vector, the generalized gate infidelity evolves roughly the same as a function of the number of IPOPT iterations. So our ability to optimize the objective function is determined mainly by the number of IPOPT iterations we are able to perform. Consequently, a method that can compute the gradient faster and therefore perform more IPOPT iterations in a fixed amount of wall time can converge to a lower objective function value. Because the higher-order methods can typically compute the gradient faster, they are able to optimize the objective function to a lower value in the same amount of time. As the target error decreases, the the higher-order methods become more effective.
@@ 5 Conclusions and Future Work
In this work, we derived a discrete adjoint method for efficiently computing gradients of objective functionals depending on the number solution of a linear system of ODEs solved by Hermite Runge-Kutta methods. Furthermore, we presented techniques for improving the computational efficiency of the method when the time-dependence of the Hamiltonian takes the typical form for systems controlled by control pulses, along with an efficient preconditioner for models where the drift Hamiltonian is diagonal.
We tested our method on the problem of designing a CNOT gate for two qudits coupled to a resonator bus, with guard levels included to ensure the validity of the model. We observed significant improvement in time and memory usage when using Hermite Runge-Kutta methods with the discrete adjoint method derived in this work (as implemented by QuantumGateDesign.jl) and comparing to the Störmer-Verlet method (as implemented by Juqbox.jl). When computing gradients to moderate levels of accuracy (relative final state error $10^{-3}$-$10^{-5}$) for the “worst-case” scenario that the system starts in its most highly excited state, we observed speedup factors of 27-143 and memory reduction factors of approximately $1,200$-$7,500$ by using high-order Hermite Runge-Kutta methods.
@@
Future work will focus on both improving the efficiency of our method, and on using the efficiency and flexibility of our method to explore new avenues in quantum optimal control. To improve the efficiency of our method, we could develop better preconditioners for solving the linear system of equations (16) at each timestep, especially for models with a non-diagonal drift Hamiltonian. We could also explore adjusting the number of timesteps and the method order to increase the accuracy as the optimization progresses, maximizing efficiency. A new application of the method could be to optimize control pulses which include the response function of the hardware used to generate them in the model.
@@ 6 Acknowledgments
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No 2235783 (SL) and by NSF under Grant Number 2436319 (DA). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. Part of this research was performed while one of the authors (SL) was visiting the Institute for Pure and Applied Mathematics (IPAM), which is supported by the National Science Foundation (Grant No. DMS-1925919). This material is based upon work supported by the U.S. Department of Energy, Office of Science, Advanced Scientific Computing Research (ASCR), under Award Number DE-SC0025424 (DA). This report was prepared as an account of work sponsored by an agency of the United States Government. Neither the United States Government nor any agency thereof, nor any of their employees, makes any warranty, express or implied, or assumes any legal liability or responsibility for the accuracy, completeness, or usefulness of any information, apparatus, product, or process disclosed, or represents that its use would not infringe privately owned rights. Reference herein to any specific commercial product, process, or service by trade name, trademark, manufacturer, or otherwise does not necessarily constitute or imply its endorsement, recommendation, or favoring by the United States Government or any agency thereof. The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof. We acknowledge Michigan State University’s support through computational resources provided by the Institute for Cyber-Enabled Research. Thanks to the creators of the Julia programming language [4] and of the packages used in the creation of this paper, including Makie.jl [10], DataFrames.jl [5], and others which do not request citations in academic literature.

### v3

@@ 4.2.2 Performing Gate Optimization
We now perform the gate optimization. We compare the performance of each order of the Hermite method for target final-time relative errors of $10^{-1}$, $10^{-3}$, $10^{-5}$, and $10^{-7}$. The optimization is managed by IPOPT (Interior Point OPTimizer) [50], a software library for solving large scale nonlinear optimization problems using a primal-dual interior point method. We configure IPOPT to perform a limited-memory quasi-Newton approximation of the Hessian of the objective function for use in line searches (second-order optimization). IPOPT also allows for box constraints on the control vector, which we use to enforce our constraint on the control pulse amplitudes.
@@
For each order and target error, we perform the optimization from ten initial control vectors, sampled from a uniform random distribution with a range of [0.005,0.005][-0.005,0.005]. The upper and lower bounds on each control parameter (enforced by IPOPT) are set to $\pm 0.05$, enforcing a maximum amplitude of 50 MHz on the B-spline envelope of each carrier wave for each control pulse, the same as the upper and lower limits on the random distribution in the stepsize experiment. The same ten initial control vectors are used for all orders and all target errors, so that every optimization starts from the same analytic control pulses. Each optimization is allowed to run for six hours before being halted.
@@
In Figure 4, for each target error we plot the generalized gate infidelity against the number of IPOPT iterations completed for each order of the Hermite method. In the figure, we do not plot the “watchdog” iterations, in which the optimizer aggressively explores new search directions, resulting in large changes in the objective function which would make the graph difficult to interpret (but the watchdog iterations are still counted in the total). These samples were collected using the same setup as the stepsize test, except we use four CPU cores instead of one and use multithreading to parallelize the forward and adjoint evolutions over the different initial and terminal conditions. In Figure 5, we plot the time evolution of the population of the computational basis states for the control found that best minimized the generalized gate infidelity for a target final state error of $10^{-7}$, which were found using the sixth-order Hermite method with 5,409 timesteps, resulting in a stepsize of $\Delta t=0.102$ nanoseconds (rounded to three significant digits).
Although the state populations in Figure 5 appear slowly varying, the dynamics of the probability amplitudes (the complex-valued components of our state vector) exhibit highly oscillatory behavior. This can be seen in Figure 6, where we plot the time evolution of the probability amplitudes for one of the initial conditions of the gate design problem when using the “best” control. The fast oscillation of the probability amplitudes accounts for the computational difficulty of numerically solving Schrödinger’s equation for this problem.
@@
Figure 4: Optimization of a CNOT gate using two qudits coupled to a resonator bus as the mode, using B-spline curves multiplied carrier waves as the control pulse ansatz, using the methods and numbers of timesteps prescribed by Table 6(a) in order to reach target relative final state errors on the orders of $10^{-1}$, $10^{-3}$, $10^{-5}$, and $10^{-7}$. We do not include the second-order method for target errors of $10^{-5}$ and $10^{-7}$ due to the memory needed to store the state history for the large number of timesteps required. The second-order method could be used with checkpointing or the “memory-lean” approach, but it is clear that at these levels of accuracy the higher-order methods are significantly faster.
@@
Figure 5: Time evolution of the state populations for the CNOT gate implemented by the best (lowest generalized infidelity) control pulses found by the optimizations performed with a target relative final state error of $10^{-7}$ for the two qudits plus resonator model (see Figure 4). The results shown here were obtained using the sixth-order Hermite method with 5,409 timesteps ($\Delta t\approx 0.102$ nanoseconds), which is the order and number of timesteps used in the optimization which found this control pulse.
Figure 6: Time evolution of the complex-valued probability amplitudes of the six states with the highest population when starting from the $|011\rangle$ state and applying the “best” control pulse for implementing a CNOT gate in the two qudits plus resonator model (see Figure 5). The results shown here were obtained using the sixth-order Hermite method with 5,409 timesteps ($\Delta t\approx 0.102$ nanoseconds), which is the order and number of timesteps used in the optimization which found this control pulse.
@@ 4.2.3 Analysis
From the plot of the mean relative error against the number of timesteps for the three subsystem example in Figure 3, we see that each order of the Hermite method exhibits the full order of convergence. From the plot of the mean gradient computation time against the number of timesteps, we can see that the computation time increases with the number of timesteps, and the asymptotic behavior of each order of the method is the same, indicating that the computation time of the higher-order methods is roughly a constant multiple of the computation time of the lower-order methods, which is to be expected. Interestingly, for the sixth through twelfth-order methods the computation time decreases around $2^{9}-2^{11}$ timesteps, for which the stepsize ranges from $1.07\geq\Delta t\geq 0.27$ nanoseconds. We attribute this to a decrease in the number of GMRES iterations required in the linear solves of the forward and adjoint evolutions due to the effectiveness of our preconditioner and our initial guess (using Taylor expansion) when $\Delta t<1$ nanosecond.
@@
The computational efficiency of the high-order methods is showcased by Tables 6(b) and 7(b). When using the initial conditions from the gate design problem and trying to reach final state errors of $10^{-3}$-$10^{-5}$, using high-order methods results in average speedup factors of approximately 33-$13$ compared to Störmer-Verlet (when choosing the most efficient order of the method for each target error). When the initial condition is the highly excited state $|933\rangle$, the average speedup factors are in the range 2727-143143!
In addition to prescribing the number of timesteps required to reach a target error, Tables 6(a) and 7(a) also highlight another advantage of the high-order methods: reduced memory requirements. The memory required to store the state history in the forward and adjoint evolutions scales linearly in the number of timesteps and the dimension of Hilbert space. Because the high-order methods can obtain highly accurate solutions with a small number of timesteps, we are often able to store the entire state history in memory. This allows us to avoid using the “memory-lean” version of the gradient calculation, which would require more computation and produce less accurate gradients.
Figure 4 highlights the effectiveness of high-order methods for the optimization process. We observe that for different order methods and different target errors, but starting from the same initial control vector, the generalized gate infidelity evolves roughly the same as a function of the number of IPOPT iterations. So our ability to optimize the objective function is determined mainly by the number of IPOPT iterations we are able to perform. Consequently, a method that can compute the gradient faster and therefore perform more IPOPT iterations in a fixed amount of wall time can converge to a lower objective function value. Because the higher-order methods can typically compute the gradient faster, they are able to optimize the objective function to a lower value in the same amount of time. As the target error decreases, the higher-order methods become more effective.
@@ 5 Conclusions and Future Work
In this work, we derived a discrete adjoint method for efficiently computing gradients of objective functionals depending on the numerical solution of a linear system of ODEs solved by Hermite Runge-Kutta methods. Furthermore, we presented techniques for improving the computational efficiency of the method when the time-dependence of the Hamiltonian takes the typical form for systems controlled by control pulses, along with an efficient preconditioner for models where the drift Hamiltonian is diagonal.
We tested our method on the problem of designing a CNOT gate for two qudits coupled to a resonator bus, with guard levels included to ensure the validity of the model. We observed significant improvement in time and memory usage when using Hermite Runge-Kutta methods with the discrete adjoint method derived in this work (as implemented by QuantumGateDesign.jl) and comparing to the Störmer-Verlet method (as implemented by Juqbox.jl). When computing gradients to moderate levels of accuracy (relative final state error $10^{-3}$-$10^{-5}$) for the “worst-case” scenario that the system starts in its most highly excited state, we observed speedup factors of 27-143 and memory reduction factors of approximately 1,2001,200-$7,500$ by using high-order Hermite Runge-Kutta methods.
@@
We also note the existence of quantum speed limits (QSL). The QSL defines the minimum evolution time it takes for a quantum system to dynamically evolve between two distinguishable quantum states [14, 6, 16]. A related concept is the minimum control time (MCT), the minimum evolution time for which the optimization of the control pulses may (in principle) be successful, according to some metric, such as the infidelity reaching a certain value [42]. In other words, given a parameterized Hamiltonian $H(t;\boldsymbol{\theta})$ and a desired gate, the gate has a minimum duration. When the coupling between qubits is weak (e.g. small cross-Kerr values) and the control pulses are amplitude-bounded due to experimental constraints, the QSL and MCT for a given gate can be large. Another advantage of our method is that high-order methods can accurately simulate over long periods of time, whereas a low-order method would be more prone to drifting from the correct solution over time. Therefore, a particularly useful application of our method would be for solving quantum optimal control problems with a large time duration imposed by the quantum speed limit or minimum control time.
@@ 6 Acknowledgments
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No 2235783 (SL) and by NSF under Grant Number 2436319 (DA). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. Part of this research was performed while one of the authors (SL) was visiting the Institute for Pure and Applied Mathematics (IPAM), which is supported by the National Science Foundation (Grant No. DMS-1925919). This material is based upon work supported by the U.S. Department of Energy, Office of Science, Advanced Scientific Computing Research (ASCR), under Award Number DE-SC0025424 (DA). This report was prepared as an account of work sponsored by an agency of the United States Government. Neither the United States Government nor any agency thereof, nor any of their employees, makes any warranty, express or implied, or assumes any legal liability or responsibility for the accuracy, completeness, or usefulness of any information, apparatus, product, or process disclosed, or represents that its use would not infringe privately owned rights. Reference herein to any specific commercial product, process, or service by trade name, trademark, manufacturer, or otherwise does not necessarily constitute or imply its endorsement, recommendation, or favoring by the United States Government or any agency thereof. The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof. We acknowledge Michigan State University’s support through computational resources provided by the Institute for Cyber-Enabled Research. Thanks to the creators of the Julia programming language [4] and of the packages used in the creation of this paper, including Makie.jl [12], DataFrames.jl [5], and others which do not explicitly request citations in academic literature.